OpenStudy (anonymous):

Evaluate the integral: e^2x-1/e^x dx

OpenStudy (ribhu):

put e^-x =t

OpenStudy (irishboy123):

if you have only just started learning this stuff, do them separately and make substitutions - which is probs what they are trying to show you. for the first part, go u = 2x, du = 2dx: 2∫exp(u) du for the second, go, u = -x, du = -dx: -∫exp(u) du in due course, you should just be doing this in your head.

OpenStudy (ribhu):

or separate the terms

OpenStudy (anonymous):

Its actually a definite integral. Can I still use u substitution for definite integrals?

OpenStudy (irishboy123):

yep

OpenStudy (irishboy123):

just make sure you cnahge the interval to one that reflects u. or convert the integral back to x and use the orig intervals.

OpenStudy (anonymous):

I just have trouble with e^x problems. Is there any way you can draw this?

OpenStudy (irishboy123):

certainly get used to drawing y = e^x, y = e^(-x) and y = ln(x). these are really really important in maths; and nature, because maths is how we model nature. but geometry/ drawing does not help IMHO. just remember this stuff: d/dx(e^x) = e^x ∫e^x dx = e^x THIS IS WHY e IS SO IMPORTANT; and this is how we calculate the value of e in the first place, ie to get this relationship. and also ∫ 1/x dx = ln(x) d/dx(ln(x) = 1/x these last 2 actually follow straight from the first but it does no harm to remember them. SO much follows from this. whether via substitution or whatever.

OpenStudy (loser66):